# Chen, Simchi-Levi, and Wang

### v1

@@
As service systems grow increasingly complex and dynamic, many interventions become localized, available and taking effect only in specific states. This paper investigates experiments with local treatments on a widely-used class of dynamic models, Markov Decision Processes (MDPs). Particularly, we focus on utilizing the local structure to improve the inference efficiency of the average treatment effect. We begin by demonstrating the efficiency of classical inference methods, including model-based estimation and temporal difference learning under a fixed policy, as well as classical A/B testing with general treatments. We then introduce a variance reduction technique that exploits the local treatment structure by sharing information for states unaffected by the treatment policy. Our new estimator effectively overcomes the variance lower bound for general treatments while matching the more stringent lower bound incorporating the local treatment structure. Furthermore, our estimator can optimally achieve a linear reduction with the number of test arms for a major part of the variance. Finally, we explore scenarios with perfect knowledge of the control arm and design estimators that further improve inference efficiency.
@@
Experimentation, A/B Testing, Markov Decision Process, Local Treatment
@@ 1 Introduction
Randomized experiments, commonly known as A/B testing, have emerged as an indispensable tool, enabling organizations across various sectors—including service, retail, manufacturing, and healthcare—to empirically test and refine their strategies for enhancing customer engagement and satisfaction (Imbens and Rubin 2015, Kohavi and Thomke 2017, Kohavi et al. 2020, Koning et al. 2022).
@@
However, as service systems become more complex and dynamic, the implementation and analysis of randomized experiments present more challenges, as well as new opportunities. One notable development in these dynamical systems is the increasing prevalence of localized treatments or policies that target specific customer interactions based on recent behaviors or other contextual factors (i.e., state). For instance, an online marketplace or service platform might offer a “welcome-back” coupon to the customer only when she has not engaged with the service for an extended period (e.g., Sephora’s Welcome Back Event Dooley 2020). Similarly, physical retail locations like pharmacies and markets may provide incentives like cash back to the customer only when she makes in-store purchases (e.g., Bed Bath & Beyond Mohammed 2015). Likewise, a typical intervention for audio stream recommender systems may involve modifying banner promotions alone, without altering other system components throughout the iterative recommendation process (e.g., Spotify Maystre et al. 2023). These localized treatments are strategic, usually aiming to optimize the lifetime value of customers by tailoring experiences and offers to individual needs and histories.
To analyze the effectiveness of local treatments in dynamic systems, we employ Markov Decision Processes (MDPs), which are commonly used for modeling complex scenarios including customer relationships (Pfeifer and Carraway 2000), inventory control (Gong and Simchi-Levi 2023), two-sided platforms (Johari et al. 2022), healthcare (Yu et al. 2021) and many others. Despite the broad applications of MDPs, experimentation on MDPs remains relatively unexplored. Existing research typically models global treatments applied uniformly across all states of an MDP. However, this approach, which ignores the potential structure of the treatment, may lead to prohibitively large variances. For example, Farias et al. (2022) demonstrate that the lower bound for any unbiased estimators can sometimes grow exponentially with the number of states. They further propose reducing the variance by sacrificing unbiasness. Other prior work, such as Shi et al. (2020), Syrgkanis and Zhan (2023) rely heavily on function approximations that may introduce additional bias, especially when the model is mis-specified. Alternatively, based on the practical observations of local treatments, in this paper we elaborate on specific treatment structures and raise the following research question:
Whether and how the local structure of treatments can be utilized to improve the inference efficiency on MDPs?
In this paper, we systematically explore the potential of leveraging the structure of local treatments to optimize data utilization from A/B testing within an MDP environment. Our findings conclusively answer this question in an affirmative and optimal manner.
1.1 An illustrative Example
@@
To be more concrete about our research question, we further introduce an application example on improving customer lifetime value, adapted from a seminal work in the marketing literature by Pfeifer and Carraway (2000) and closely related to customer retention and churn management (Lemmens and Gupta 2020, Kanoria et al. 2024). This example will be revisited throughout the paper.
@@
Drawing from the classic Recency, Frequency, Monetary value (RFM) framework, it is often posited by companies that the probability a customer will purchase at the end of any period is solely a function of her recency—the number of periods since her last purchase. Pfeifer and Carraway (2000) model recency as the state space of a MDP and characterizes customer lifetime value using value functions. The state transitions occur as follows: each purchase resets the customer’s recency to one; conversely, if no purchase is made, the customer’s recency increases by one period. When a customer’s recency is high (for example, $\geq$5), Pfeifer and Carraway (2000) aggregate these instances into a single state, denoted as $s^{5}$, and regard it as representing an inactive customer.
@@ 1.2 Overview of Main Results
Efficient Inference under a Fixed Policy. Instead of simultaneously evaluating both the treatment and control policies, we initially focus on presenting estimation methods for the long-term cumulative rewards of a single, specific policy, say, either the treatment or the control.
The two principal approaches used in the current literature are model-based estimation and temporal difference (TD) learning. Model-based estimation involves directly estimating all parameters of the MDP under a fixed policy to subsequently compute long-term rewards. Conversely, TD learning is more straightforwards to fit the long-term rewards according to the Bellman equation. Despite their distinct approaches, we establish that both methods are equivalent and demonstrate asymptotic normality. This means that the two estimators yield identical results and share confidence intervals. This equivalence allows the theoretical results derived from one method to be applicable to the other, providing the flexibility to select the method that is simpler to analyze. Moreover, we derive a variance lower bound for all unbiased estimators using a constrained Cramér-Rao bound. This highlights the inherent challenges of the inference task and confirms that both methods are asymptotically efficient compared to all other unbiased estimators, which may be of independent interest for the reinforcement learning (RL) literature.
@@
Having demonstrated the variance reduction achieved through SST, we also expand our framework and generalize the results to encompass more general local treatments. This extension implies that the new treatment may be applicable to multiple states, broadening the scope and applicability of our results.
Moreover, we reveal that under SST, the treatment effect remains consistently positive (or negative) for all states before and after the intervention. This indicates that if the new treatment benefits one state, it benefits all others as well. Also, we discuss that our information sharing concept is orthogonal to and compatible with the bias-variance trade-offs (e.g., Farias et al. 2022) and function approximation approaches (e.g., Shi et al. 2020, Syrgkanis and Zhan 2023) developed for the global treatments. This means that many developments under the global treatments can be further improved with the presence of the local treatments by our information sharing idea.
@@ 1.3 Related Works
Structure-aware Experiment Design. Experimental design originated in economics with basically model-free methods. In recent years, researchers in operations research and management science instead realize that there are many useful structures in operations scenarios to understand, analyze and improve experimentation. Bojinov et al. (2023), Hu and Wager (2022), Ni et al. (2023), Jia et al. (2023), Xiong et al. (2024) consider the switchback design under various scenarios, from temporal interference to spatial interference, from stationary environment to non-stationary environment, and from theory driven perspective to data driven perspective. Glynn et al. (2020), Munro et al. (2024), Wager and Kuang (2021), Bajari et al. (2021), Johari et al. (2022), Zhu et al. (2024) study the experiments happening on two-sided markets and a common practice is to model the system as a continuous-time Markov chain. Li et al. (2015) leverage the sparsity structure of cross-product demand elasticities to greatly reduce the number of experiments required. Li et al. (2023) carefully take the queuing structure into account and propose efficient estimators under the stochastic congestion. Xiong et al. (2023) investigate the setting with multiple units where the starting time of the treatment may vary by unit. Zhan et al. (2021, 2023) study the post-experiment inference for contextual bandits, which does not capture the state transition as our MDPs. Simester et al. (2020) pioneer at considering treatment structures and suggest comparing outcomes only for customers exposed to different treatment and control actions to improve efficiency empirically. Compared to theirs, we further incorporate complex transition dynamics and provide theoretical guarantees of estimations. The long-term effect with the long-term treatment under the short-term measurement is recently studied by Huang et al. (2023). This paradigm of structure-aware experiment design has witnessed great success in theory and practice in above works with either significant reduction of bias or substantial improvement of efficiency.
@@ 2 Model
Consider a (non-episodic, finite-state, infinite-horizon) MDP with state space ${\mathcal{S}}:=\{s^{i}\}_{i\in[K]}$ with cardinality $K$. At time $t\in\mathbb{N}$, we denote the state of the MDP by $s_{t}\in{\mathcal{S}}$. Associated with each state is a set of available actions, ${\mathcal{A}}$, which influence the state transitions through the transition probability function $P:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\rightarrow[0,1]$. Here, the next state $s_{t+1}$ follows the distribution $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$.
For each state-action pair $(s,a)$, the instant reward $r$ earned is independently generated by a Gaussian distribution [^1] With standard arguments, our results can be extended to the setting where the second moment of the noise is finite., i.e., $r\sim{\mathcal{N}}(r(s,a),\sigma_{r}^{2}(s,a))$, where we assume $\sigma^{2}_{r}(s,a)<\infty$ for all $(s,a)$ pairs.
A policy $\pi:{\mathcal{S}}\rightarrow{\mathcal{A}}$ maps states to (random) actions. The value function is defined as the cumulative discounted reward starting from state $s\in{\mathcal{S}}$ under a policy $\pi$,
@@
We then introduce a special class of treatment policies that we call the Single-State Treatment (SST). Specifically, there exists a pre-determined “crucial” state, which without loss of generality, we assume to be $s^{1}$. The treatment of interest is applied exclusively when the system’s state, $s_{t}$, coincides with $s^{1}$.
Formally, we assume that ${\mathcal{A}}=\{{\mathfrak{c}},{\mathfrak{t}}\}$ regardless of states, while the action ${\mathfrak{t}}$, representing the new treatment, is only applicable in the crucial state $s^{1}$, while ${\mathfrak{c}}$ denotes the control or default action. This restriction ensures that the treatment is targeted and its effects are isolated to the crucial state.
@@ 3 Preliminaries: Inference under A Fixed Policy
It is worth noting that any estimation $\widehat{\bm{P}}$ is always a transition matrix. Thus the matrix $({\bm{I}}-\gamma\widehat{\bm{P}})$ is always non-singular. Therefore one can verify that the unique solution to Eq. (4) is
$\widehat{\bm{V}}^{\textnormal{TD}}=({\bm{I}}-\gamma\widehat{\bm{P}})^{-1}\widehat{\bm{r}}\,,$
@@
13:  $\widehat{P}(s^{\prime}\mid s,a)=N(s,a,s^{\prime})/(N(s,a)\vee 1)$, $\forall s,s^{\prime}\in{\mathcal{S}},a\in{\mathcal{A}}$
@@
10:  The obtained trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$; Discount factor $\gamma$.
21:  Solve the following optimization problem Eq. (3),
@@ 3.1 Asymptotic Inference of Value Function
Recall that $\rho^{\pi}_{s^{1}}(s)$ is the discounted state-visitation measure defined in Definition 2.1. The asymptotic variance of estimating $V(s^{1})$ is influenced not only by the state $s^{1}$ itself but also by all the other states. This is consistent with our understanding that our objective is the long-term value and the samples are drawn from a MDP. Although the variance depends on all states, we want to specially highlight that $\sigma^{2}_{\mathrm{MB}}(s^{1})$ can be decomposed into $K$ distinct terms, where each term corresponds almost exclusively to one state. Furthermore, each term is inversely proportion to the stationary probability $\mu^{\pi}(s)$. This implies that states which are less frequently visited pose greater challenges in sampling, potentially leading to increased variance. Additionally, each term has two parts, $\sigma_{\pi}^{2}(s)$ and $\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}$, which correspond to the (short-term) variance of the instant rewards and the (long-term) variance related to transitions to the other states.
@@ 5.2 Lower Bound for Experimentation with SST
Consider a MDP with state space ${\mathcal{S}}\coloneqq\left\{{s^{i}}\right\}_{i\in[K]}$, where $s^{1}$ is still assumed to be the crucial state of treatment. Unlike the previous model in Section 2 with two test arms, the action set ${\mathcal{A}}=\left\{{{\mathfrak{t}}^{j}}\right\}_{j\in[n]}$ now includes $n$ possible actions. Without loss of generality, we let ${\mathfrak{t}}^{1}$ denote the control action. Recall Example 1.1, a plausible experiment might involve the company testing multiple coupon values to determine which one maximizes customer lifetime value. This implies $n$ SST polices $\Pi\coloneqq\left\{{\pi^{i}}\right\}_{i\in[n]}$ with each $\pi^{i}$ satisfying: $\pi^{i}(s^{1})\equiv{\mathfrak{t}}^{i}$ (i.e., $\pi^{i}({\mathfrak{t}}^{i}\mid s^{1})=1$) and $\pi^{i}(s)\equiv{\mathfrak{t}}^{1}$ (i.e., $\pi^{i}({\mathfrak{t}}^{1}\mid s)=1$) for all $s\in{\mathcal{S}}\setminus s^{1}$. Additionally, each SST policy is associated with its own transition matrix ${\bm{P}}^{{\mathfrak{t}}^{i}}$ and mean reward ${\bm{r}}^{{\mathfrak{t}}^{i}}$.

### v2

@@
Utilizing randomized experiments to evaluate the effect of short-term treatments on the short-term outcomes has been well understood and become the golden standard in industrial practice. However, as service systems become increasingly dynamical and personalized, much focus is shifting toward maximizing long-term cumulative outcomes, such as customer lifetime value, through lifetime exposure to interventions. To bridge this gap, we investigate the randomized experiments within dynamical systems modeled as Markov Decision Processes (MDPs). Our goal is to assess the impact of treatment and control policies on long-term cumulative rewards from relatively short-term observations. We first develop optimal inference techniques for assessing the effects of general treatment patterns. Furthermore, recognizing that many real-world treatments tend to be fine-grained and localized for practical efficiency and operational convenience, we then propose methods to harness this localized structure by sharing information on the non-targeted states. Our new estimator effectively overcomes the variance lower bound for general treatments while matching the more stringent lower bound incorporating the local treatment structure. Furthermore, our estimator can optimally achieve a linear reduction with the number of test arms for a major part of the variance. Finally, we explore scenarios with perfect knowledge of the control arm and design estimators that further improve inference efficiency.
@@
Experimentation, A/B Testing, Lifetime Value, Markov Decision Process, Local Treatment
@@ 1 Introduction
Randomized experiments, commonly known as A/B testing, have emerged as an indispensable tool, enabling organizations across various sectors—including service, retail, manufacturing, and healthcare—to empirically test and refine their strategies for enhancing customer engagement and satisfaction (Imbens and Rubin 2015, Kohavi and Thomke 2017, Kohavi et al. 2020, Koning et al. 2022). In particular, A/B testing is well-established and widely used to compare short-term or immediate outcomes (e.g., whether a customer clicks on an advertisement) resulting from short-term treatments (e.g., displaying a new version of the advertisement).
However, as service systems grow more complex and dynamic, the focus is increasingly shifting from immediate or short-term outcomes to long-term effects on customers or users who are continually exposed to new interventions. For example, online retailers may employ strategies such as regular coupon distribution or cash back, aiming to trade-off short-term profit reductions for potential long-term revenue gains (Simester et al. 2006). Similarly, user interface updates on web platforms tend to have continual exposure to customers after their launches, thus impacting user experience and engagement over extended periods (Huang et al. 2023, Maystre et al. 2023). These practical scenarios require estimating the impact of long-term treatments on long-term cumulative outcomes, such as customer lifetime value, using relatively short-term observations from randomized experiments.
@@
In this work, we investigate the randomized experiments on Markov Decision Processes (MDPs), which are widely used to model dynamic service systems over extended time horizons including customer relationships (Pfeifer and Carraway 2000), inventory control (Gong and Simchi-Levi 2023), two-sided platforms (Johari et al. 2022), mobile health (Zhang et al. 2023) and many others. Additionally, as modern systems become increasingly complex with numerous states, the practical challenges of implementing entirely new global policies intensify. Such policies could be overly aggressive, in terms of posing significant risks and requiring substantial infrastructure changes. Consequently, many treatments in practice are becoming more fine-grained and localized. This means that the new treatments may only apply on a small number of targeted states locally. For instance, in the customer relationships, the customer’s states are referring to her recent behaviors or other contextual factors. An online marketplace or service platform might offer a “welcome-back” coupon to the customer only when she has not engaged with the service for an extended period (e.g., Sephora’s Welcome Back Event Dooley 2020). Similarly, physical retail locations like pharmacies and markets may provide incentives like cash back to the customer only when she makes in-store purchases (e.g., Bed Bath & Beyond Mohammed 2015). Likewise, a typical intervention for audio stream recommender systems may involve modifying banner promotions alone, without altering other system components throughout the iterative recommendation process (e.g., Spotify Maystre et al. 2023). These localized treatments are strategic, usually aiming to optimize the lifetime value of customers by tailoring experiences and offers to individual needs and histories. Given the prevalence of local treatments in practice, our research goal is to explore the potential of leveraging local treatment structure to optimize data utilization for efficient experimentation on the long-term cumulative reward within an MDP environment.
1.1 An illustrative Example On Customer Lifetime Value
@@
To be more concrete about our research question, we further introduce an application example on improving customer lifetime value, adapted from a seminal work in the marketing literature by Pfeifer and Carraway (2000) and closely related to customer retention and churn management (Hamilton and Singal 2023, Lemmens and Gupta 2020, Kanoria et al. 2024). This example will be revisited throughout the paper.
@@
Drawing from the classic Recency, Frequency, Monetary value (RFM) framework, it is often posited by companies that the probability a customer will purchase at the end of any period is solely a function of her recency—the number of periods since her last purchase. Pfeifer and Carraway (2000) model recency as the state space of an MDP and characterize customer lifetime value via the value function, which is the discounted sum of revenue collected in the long-run.The state transitions occur as follows: each purchase resets the customer’s recency to one; conversely, if no purchase is made, the customer’s recency increases by one period. When a customer’s recency is high (for example, $\geq$5), Pfeifer and Carraway (2000) aggregate these instances into a single state, denoted as $s^{5}$, and regard it as representing an inactive customer.
@@ 1.2 Overview of Main Results
Efficient Inference under a Fixed Policy. Instead of simultaneously evaluating both the treatment and control policies, we initially focus on presenting estimation methods for the long-term cumulative rewards of a single, specific policy, say, either the treatment or the control. The two principal approaches used in the current literature are model-based estimation and temporal difference (TD) learning. Model-based estimation involves directly estimating all parameters of the MDP under a fixed policy to subsequently compute long-term rewards. Conversely, TD learning is more straightforwards to fit the long-term rewards according to the Bellman equation. Despite their distinct approaches, we establish that both methods are equivalent and demonstrate asymptotic normality. This means that the two estimators yield identical results and share confidence intervals. This equivalence allows the theoretical results derived from one method to be applicable to the other, providing the flexibility to select the method that is simpler to analyze. Moreover, we derive a variance lower bound for all unbiased estimators using a constrained Cramér-Rao bound. This highlights the inherent challenges of the inference task and confirms that both methods are asymptotically efficient compared to all other unbiased estimators, which may be of independent interest for the reinforcement learning (RL) literature.
@@
Having demonstrated the variance reduction achieved through SST, we also expand our framework and generalize the results to encompass more general local treatments. This extension implies that the new treatment may be applicable to multiple states, broadening the scope and applicability of our results. We also test the effectiveness of our design on a real-world Sepsis simulator. Moreover, we reveal that under SST, the treatment effect remains consistently positive (or negative) for all states before and after the intervention. This indicates that if the new treatment benefits one state, it benefits all others as well. Also, we discuss that our information sharing concept is orthogonal to and compatible with the bias-variance trade-offs (e.g., Farias et al. 2022) and function approximation approaches (e.g., Shi et al. 2020, Syrgkanis and Zhan 2023) developed for the global treatments. This means that many developments under the global treatments can be further improved with the presence of the local treatments by our information sharing idea.
@@ 1.3 Related Works
Structure-aware Experiment Design. Experimental design originated in economics with basically model-free methods. In recent years, researchers in operations research and management science instead realize that there are many useful structures in operations scenarios to understand, analyze and improve experimentation. Bojinov et al. (2023), Hu and Wager (2022), Ni et al. (2023), Jia et al. (2023), Xiong et al. (2024) consider the switchback design under various scenarios, from temporal interference to spatial interference, from stationary environment to non-stationary environment, and from theory driven perspective to data driven perspective. Glynn et al. (2020), Munro et al. (2024), Wager and Kuang (2021), Bajari et al. (2021), Johari et al. (2022), Zhu et al. (2024), Weng et al. (2024) study the experiments happening on two-sided markets and a common practice is to model the system as a continuous-time Markov chain. Li et al. (2015) leverage the sparsity structure of cross-product demand elasticities to greatly reduce the number of experiments required. Li et al. (2023) carefully take the queuing structure into account and propose efficient estimators under the stochastic congestion. Xiong et al. (2023) investigate multi-unit settings with varying treatment starting times. Zhan et al. (2021, 2023) study the post-experiment inference for contextual bandits, which does not capture the state transition as our MDPs. Simester et al. (2020) pioneer at considering treatment structures and suggest comparing outcomes only for customers exposed to different treatment and control actions to improve efficiency empirically. Compared to theirs, we further incorporate complex transition dynamics and provide theoretical guarantees of estimations. The long-term effect with the long-term treatment under the short-term measurement is recently studied by Huang et al. (2023). This paradigm of structure-aware experiment design has witnessed great success in theory and practice in above works with either significant reduction of bias or substantial improvement of efficiency.
@@ 2 Model
Consider a (non-episodic, finite-state, infinite-horizon) MDP with state space ${\mathcal{S}}:=\{s^{i}\}_{i\in[K]}$ with cardinality $K$. At time $t\in\mathbb{N}$, we denote the state of the MDP by $s_{t}\in{\mathcal{S}}$. Associated with each state is a set of available actions, ${\mathcal{A}}$, which influence the state transitions through the transition probability function $P:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\rightarrow[0,1]$. Here, the next state $s_{t+1}$ follows the distribution $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$. For each state-action pair $(s,a)$, the instant reward $r$ earned is independently generated by a Gaussian distribution [^1] With standard arguments, our results can be extended to the setting where the second moment of the noise is finite., i.e., $r\sim{\mathcal{N}}(r(s,a),\sigma_{r}^{2}(s,a))$, where we assume $\sigma^{2}_{r}(s,a)<\infty$ for all $(s,a)$ pairs. A policy $\pi:{\mathcal{S}}\rightarrow{\mathcal{A}}$ maps states to (random) actions. The value function is defined as the cumulative discounted reward starting from state $s\in{\mathcal{S}}$ under a policy $\pi$,
@@
We then introduce a special class of treatment policies that we call the Single-State Treatment (SST). Specifically, there exists a pre-determined “crucial” state, which without loss of generality, we assume to be $s^{1}$. The treatment of interest is applied exclusively when the system’s state, $s_{t}$, coincides with $s^{1}$. Formally, we assume that ${\mathcal{A}}=\{{\mathfrak{c}},{\mathfrak{t}}\}$ regardless of states, while the action ${\mathfrak{t}}$, representing the new treatment, is only applicable in the crucial state $s^{1}$, while ${\mathfrak{c}}$ denotes the control or default action. This restriction ensures that the treatment is targeted and its effects are isolated to the crucial state.
@@ 3 Preliminaries: Inference under A Fixed Policy
3:  $\widehat{P}(s^{\prime}\mid s,a)=N(s,a,s^{\prime})/(N(s,a)\vee 1)$, $\forall s,s^{\prime}\in{\mathcal{S}},a\in{\mathcal{A}}$
@@
1:  Solve the following optimization problem Eq. (3),
@@ 3.1 Asymptotic Inference of Value Function
Recall that $\rho^{\pi}_{s^{1}}(s)$ is the discounted state-visitation measure defined in Definition 2.1. The asymptotic variance of estimating $V(s^{1})$ is influenced not only by the state $s^{1}$ itself but also by all the other states. This is consistent with our understanding that our objective is the long-term value and the samples are drawn from an MDP. Although the variance depends on all states, we want to specially highlight that $\sigma^{2}_{\mathrm{MB}}(s^{1})$ can be decomposed into $K$ distinct terms, where each term corresponds almost exclusively to one state. Furthermore, each term is inversely proportion to the stationary probability $\mu^{\pi}(s)$. This implies that states which are less frequently visited pose greater challenges in sampling, potentially leading to increased variance. Additionally, each term has two parts, $\sigma_{\pi}^{2}(s)$ and $\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}$, which correspond to the (short-term) variance of the instant rewards and the (long-term) variance related to transitions to the other states.
@@ 5.2 Lower Bound for Experimentation with SST
Consider an MDP with state space ${\mathcal{S}}\coloneqq\left\{{s^{i}}\right\}_{i\in[K]}$, where $s^{1}$ is still assumed to be the crucial state of treatment. Unlike the previous model in Section 2 with two test arms, the action set ${\mathcal{A}}=\left\{{{\mathfrak{t}}^{j}}\right\}_{j\in[n]}$ now includes $n$ possible actions. Without loss of generality, we let ${\mathfrak{t}}^{1}$ denote the control action. Recall Example 1.1, a plausible experiment might involve the company testing multiple coupon values to determine which one maximizes customer lifetime value. This implies $n$ SST polices $\Pi\coloneqq\left\{{\pi^{i}}\right\}_{i\in[n]}$ with each $\pi^{i}$ satisfying: $\pi^{i}(s^{1})\equiv{\mathfrak{t}}^{i}$ (i.e., $\pi^{i}({\mathfrak{t}}^{i}\mid s^{1})=1$) and $\pi^{i}(s)\equiv{\mathfrak{t}}^{1}$ (i.e., $\pi^{i}({\mathfrak{t}}^{1}\mid s)=1$) for all $s\in{\mathcal{S}}\setminus s^{1}$. Additionally, each SST policy is associated with its own transition matrix ${\bm{P}}^{{\mathfrak{t}}^{i}}$ and mean reward ${\bm{r}}^{{\mathfrak{t}}^{i}}$.